OpenStudy (anonymous):

In the field of thermodynamics, Rankine is the unit used to measure temperature. One can convert temperature from Celsius into Rankine using the formula C =R/1.8-273 , What is the temperature in Celsius corresponding to R degrees Rankine?

OpenStudy (anonymous):

9/5(C-273) 9/5(C+273) 5/9(C-273) 5/9(c+273)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@pooja195

OpenStudy (rational):

C =R/1.8-273 i think they want you solve R

OpenStudy (anonymous):

ok

OpenStudy (rational):

do you know how to solve R in above equation ?

OpenStudy (anonymous):

I dont

OpenStudy (anonymous):

would the 273 add or something?

OpenStudy (rational):

Yes! start by adding 273 both sides: C+273 =R/1.8

OpenStudy (rational):

next what you wanto do ?

OpenStudy (anonymous):

something with that 1.8

OpenStudy (rational):

yes get rid of it

OpenStudy (anonymous):

have no clue on that part

OpenStudy (rational):

multiply 1.8 both sides

OpenStudy (rational):

you get 1.8(C+273) =R

OpenStudy (rational):

notice that 1.8 = 18/10 = 9/5

OpenStudy (anonymous):

9/5(c+273)

OpenStudy (rational):

looks good!

OpenStudy (anonymous):

thx bro

OpenStudy (rational):

yw